OpenStudy (anonymous):

Find f (g (x)) for f (x) = 6x + 3 and g(x) = 5x – 2.

OpenStudy (anonymous):

plug in g(x) for every x in f(x) it's like finding f(3) but you're finding f(g(x)) f(g(x)) = 6(g(x)) + 3 = 6(5x -2) + 3

OpenStudy (anonymous):

f(5x-2)=30x-9??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks :)